# Large Language Model Strategic Reasoning Evaluation through Behavioral Game Theory

### v2

@@ 5.2 Baseline Evaluation
Table 2: Strategic Reasoning Performance ($\tau$) Across Different Game Settings.
@@
Figure 2: Performance of all models across game types
Table 3: Strategic Reasoning Performance ($\tau$) Across Different Game Settings in CoT.
@@
The baseline setup provides a foundational understanding of how LLMs perform in strategic games when relying solely on their inherent reasoning mechanisms. The results are presented in Table 2. The distribution of $\tau$ is shown in Figure 2 for all models.
Among all evaluated models, GPT-o1, GPT-o3-mini, and DeepSeek-R1 outperform others, with GPT-o1 ranking first in 4 games, GPT-o3-mini leading in 7 games, and DeepSeek-R1 securing the top position in 5 games. Some games have co-leaders, where multiple models achieve the highest strategic reasoning level. In these cases, certain LLMs consistently reach the optimal strategy—often aligning with Nash Equilibrium (NE)—demonstrating their ability to fully reason through the game structure. However, not all models achieve this level of performance, highlighting variations in reasoning depth. This underscores both the strengths of high-performing models in structured settings and the effectiveness of our framework in distinguishing optimal decision-making from suboptimal or inconsistent reasoning. Additionally, our findings reveal significant variations in reasoning capabilities across different game types. GPT-o1 consistently ranks among the top models in competitive and incomplete-information games, suggesting an optimization for rational, goal-oriented decision-making and adversarial reasoning. DeepSeek-R1 excels in cooperative and mixed-motive games, likely due to its enhanced ability to model social interactions and optimize mutual benefits under reinforcement learning rewards. GPT-o3-mini, meanwhile, demonstrates strong performance across a broader range of games, particularly excelling in cooperative, mixed-motive, and incomplete-information games. Its success in both social and strategic reasoning tasks suggests a balanced capability in handling trade-offs, adapting to uncertainty, and leveraging probabilistic inference in decision-making.
The performance gap between stronger models, such as DeepSeek-R1, GPT-o1 and GPT-o3-mini, and weaker ones, like DeepSeek V2.5, DeepSeek V3 and GPT-4o, is notable across most games. Nevertheless, advanced models do not always dominate all game types. Smaller models like Gemma-2-27B and LLaMA-3.1-8B sometimes match or outperform larger counterparts, and DeepSeek-R1 is a top model despite not being the largest. Additionally, R1-distilled models can outperform significantly larger models, despite being derived from a much smaller base model. This suggests that model size is not the sole factor in reasoning quality. Instead, context, model structure, and training data likely play a crucial role, highlighting the strong influence of contextual adaptation on LLMs’ reasoning abilities.
Comparing with vs. without Chain-of-Thought
The performances of the strategic reasoning level after implementing the CoT prompt are presented in Table 3. Intuitively, many believe CoT can likely enhance a model’s reasoning ability. However, our results show that CoT’s impact on strategic reasoning is nuanced, with no consistent improvement across all models and game types.
Figure 3: Comparison of Tau between baseline and CoT in the Competitive-Base game
Figure 3 shows an example comparing $\tau$ of each model in the Competitive-Base game with and without CoT prompting. While some models improve, others remain unchanged or even decline. Notably, a model that gains from CoT in one game may not in another. This variation may reflect how each model processes and benefits from step-by-step explicit reasoning—some are better at translating CoT text into genuinely more iterative strategies, whereas others might get sidetracked or produce inconsistent lines of thought. This is consistent with prior findings suggesting that the "CoT does not always help" [26, 13], particularly in strategic settings. Nonetheless, a close examination of Figure 4 reveals that the bars are aggregated around $\tau<1.8$, meaning more models have improved $\tau$ when original reasoning capability is low, indicating the potential of the CoT prompting technique to enhance reasoning levels. That is, when a model’s reasoning capability is very low, CoT can provide a slight improvement, elevating it to a moderate level.
Figure 4: Comparison of the distribution of Tau between baseline and CoT prompting for all models across all game types
For those low-reasoning capability systems, CoT is more likely to provide enough structured guidance to improve their reasoning process, leading to modest but meaningful increases in $\tau$. This suggests that CoT can partially compensate for weaker internal representations of strategic play, giving the model a chance to articulate an “opponent’s perspective” or weigh payoffs more carefully. In contrast, as shown in Figure 3, stronger models, such as GPT-4o, which already demonstrate moderate $\tau$ in baseline evaluation, do not always see significant additional gains from CoT. One possibility is that these models internally perform iterative reasoning even without explicit prompts, so the CoT text adds less incremental benefit to their final outputs.
Comparison with Human
We compare all model settings, incorporating both internal and external CoT where applicable, against human subjects in the same game setting [70, 67]. Several key observations emerge, as shown in Figure 5. A total of 17 models outperform humans in this benchmark, underscoring how certain LLMs can excel beyond typical human reasoning levels in specific strategic scenarios. Models such as GPT-4 CoT, DSk-R1_Qwen-14B, and other large-parameter systems often occupy top ranks, suggesting that a combination of scale, refined architectures, and careful prompting (e.g., CoT) can yield substantial gains in strategic reasoning.
Despite being outperformed by the strongest models, humans rank in the middle tier, surpassing multiple LLMs. This suggests that while SOTA models exhibit advanced strategic reasoning, human decision-making remains robust and sophistication. It is important to note that this evaluation represents just one specific game setting, providing a snapshot rather than a definitive ranking across all strategic tasks. As performance varies across different types of games, a broader analysis is necessary to fully understand the comparative strengths and weaknesses of each model in diverse strategic environments.
Figure 5: Comparison among human participants and LLMs on the SW10 game [70]
While the reasoning levels demonstrated by the models provide valuable insights, it is important to question whether "higher is always better." In fact, alignment with human expectations and contextual requirements is paramount. Some models outperform humans in structured scenarios, but such dominance does not necessarily translate into broader applicability or ethical alignment with human decision-making norms. For example, in scenarios requiring fairness, trust, or cooperation—such as resource allocation or collaborative decision-making—models that align closely with human reasoning may be preferable. Conversely, outperforming human reasoning may be beneficial in structured, high-stakes tasks requiring precise optimization. Ultimately, reasoning capabilities must be calibrated to respect the objectives and constraints of each scenario. While surpassing humans demonstrates technical sophistication, achieving alignment with human values and societal goals remains the more meaningful and complex challenge.
5.3 With Encoded Demographic Feature Evaluation
After establishing each model’s strategic reasoning level, we explore and evaluate the correlation between demographic variables and the model’s reasoning capability through Ordinary Least Square (OLS) regression. Specifically, we estimate the following regression model:
$\tau_{m}=\alpha+\beta D+\epsilon$
where $\tau_{m}$ represents the estimated reasoning level of model $m$, $D$ is a vector of demographic variables (e.g., gender, age, education, and political affiliation) and $\epsilon$ is the error term. The coefficient $\beta$ and its associated significant level captures the effect of demographic attributes on the model’s reasoning performance.
Several models exhibited distinct levels of “sensitivity” to these variables, as shown in Appendix Table 4.
Figure 6: Selected average parameter values where regression coefficients (indicators of the relationship between variables in a statistical model) are significant in this category with non-external-CoT models.
GPT-4o and Gemma-2 illustrate these effects as notable cases. For instance, in Figure 6(a), GPT-4o’s reasoning levels varied by gender, with female features increasing and male features decreasing reasoning levels. Similarly, Figure 6(b) illustrates how Gemma-2 exhibited pronounced differences based on sexual orientation, with reasoning levels spiking for heterosexuals compared to other minority categories. These findings suggest that the model’s internal processing may partially rely on linguistic cues tied to users’ personal backgrounds.
By contrast, LLaMA-3.1-8B and QwQ-32B showed fewer demographic correlations, indicating less sensitivity to user attributes in strategic depth allocation. Some models, such as LLaMA-3.1-405B, fell into an intermediate range, with only a handful of demographic variables—and often ones related to sexual orientation or political orientation—reaching standard or near-standard significance. These findings may indicate which types of demographic features are likely to trigger bias within LLMs. Although these results do not necessarily point to large-scale biases, they do indicate that certain demographic signals can shape how these models reason about or respond to strategic prompts, potentially raising fairness concerns.
We then introduced CoT prompting to see whether explicit step-by-step reasoning heightened or diminished the observed correlations and presented the results in Appendix Table 5. Overall, more demographic variables became significant under CoT—even for models that were previously neutral—reflecting a tendency for CoT to alter how the model weighs user cues. For example, Figure 7 shows that Gemma-2 displayed marked increases for "Jewish" features, while InternLM, initially neutral, developed religious disparities. These findings highlight fairness concerns, as CoT can both expose latent biases and introduce new correlations, making fairness dynamic across prompts. CoT’s ability to introduce new correlations underscores that fairness is not static across prompt variations. If a model prioritizes certain demographic signals, CoT may amplify rather than mitigate bias. This underscores the need for ongoing, prompt-specific bias audits as its use expands.
Biases in distilled models present a notable challenge. While distillation improves the reasoning ability of smaller models, this enhancement is most evident in cooperation and mixed-motive game settings. However, it also introduces additional biases, which must be carefully considered. In the vanilla form of LLaMA, even after embedding CoT, bias remains limited, with only one or two features affecting reasoning levels. However, knowledge distillation introduces additional biases in both LLaMA-8B and LLaMA-70B. Compared to CoT alone, distillation amplifies sensitivity to demographic factors, as DSk-R1_LLaMA70B exhibits a significant bias toward age and race. These findings highlight the importance of carefully monitoring knowledge distillation. If malicious data are used to train the teacher model, the resulting biases may propagate to the student model, potentially leading to harmful behaviors. More critically, even with benign training data, hidden biases can transfer during distillation, underscoring the need for caution when applying this technique.
6 Conclusion and Discussion
This study sets a new milestone in LLM strategic reasoning evaluation by assessing strategic capability while accounting for contextual influences, addressing the limitations of benchmarks focused solely on NE. We find that models that perform well on established benchmarks of text, math, etc., do not necessarily always have higher reasoning levels than others. Our evaluation also uncovers CoT’s limitations, showing that its benefits are not universal and can sometimes degrade reasoning performance. This insight contributes to ongoing discussions on enhancing LLM reasoning workflows, particularly in complex decision-making scenarios. Finally, our analysis of demographic sensitivity highlights an often-overlooked aspect of strategic reasoning—how biases can emerge even in high-performing models. Furthermore, we find that while model distillation can enhance the reasoning ability of smaller models in specific game types, it also introduces additional bias toward demographic features.
Beyond Scale and Benchmark Performance. While larger models generally perform well, DeepSeek-R1 achieves competitive results despite its smaller scale. Its success, driven by reinforcement learning and fine-tuned optimization [36], underscores the importance of training methodologies in shaping reasoning abilities.
Although our approach provides a comprehensive evaluation of multi-faceted decision-making, further research is needed to clarify how factors such as mathematical and logical proficiency, post-training techniques, and model scale influence reasoning. These relationships remain unclear, but our work establishes an initial link. Building reliable AI agents requires understanding these dependencies, as improvements in isolated tasks do not necessarily enhance overall reasoning. Identifying and optimizing key hyperparameters is crucial to achieving robust reasoning capabilities.
CoT is Not Uniformly Effective. While CoT prompting enhances reasoning in language models, its effectiveness depends on task comprehension and alignment with the intended goal. Our analysis reveals that when properly aligned, CoT scaffolds structured logic and improves performance; however, misaligned CoT prompts can amplify errors, akin to optimizing a flawed policy in reinforcement learning. This challenges the proposal that human-designed CoT is universally optimal for step-by-step reasoning, as LLMs may rely on latent reasoning mechanisms—such as parallel or hierarchical logic—that deviate from strict sequential processing. Empirical evidence from DeepSeek-R1 supports this view, as it successfully achieves SOTA performance through reinforcement learning, integrating CoT-like reasoning without enforcing explicit sequential structures. Instead, it dynamically adapts reasoning pathways through reward-driven optimization. Unlike static rule-based systems or predefined sequential reasoning processes, RL-based models iteratively refine their strategies based on feedback from the environment. This allows AI agents to develop flexible, context-aware reasoning mechanisms rather than relying solely on fixed logical structures. These findings highlight a fundamental open question: how to align external prompting strategies with a model’s intrinsic reasoning processes, a challenge crucial for advancing robust and interpretable AI systems.
Fairness and Bias in AI Reasoning. Superior strategic reasoning does not equate to ethical decision-making. When demographic cues are introduced, models adjust their strategies based on the given role, indicating that AI-driven decisions can encode and amplify societal biases. This is especially concerning in multi-agent interactions, where prioritizing strategic efficiency alone may lead to inequitable or adversarial outcomes.
In interactive settings, LLMs that adjust strategies based on demographics risk perpetuating bias in negotiations and resource allocation, reinforcing real-world inequalities over time. This presents a critical challenge to balance strategic reasoning with fairness. Unconstrained optimization may encourage exploitative tactics, while rigid fairness constraints could hinder adaptability. Future development must move beyond performance metrics to integrate ethical safeguards, ensuring that LLMs engage in fair and justifiable decision-making in dynamic multi-agent environments.
Limitations: While we examine CoT prompting and links between individual and comprehensive benchmarks, the precise causal chain between prompts and final decision-making remains unexplored. Future research could focus on digging into how individual skills, such as logical reasoning, mathematical ability, and contextual understanding, influence an LLM’s overall reasoning capabilities. Additionally, our evaluation framework is semi-dynamic, relying on non-real-time multi-agent scenarios. In real-time interactions, the same methodology can also apply, but results may differ due to dynamic rectifications. Such settings may reveal different reasoning levels or widen performance gaps between top-tier and lower-tier models due to the interactive nature of these evaluations. Future work may explore these scenarios for deeper insights into LLM behavior.
Appendices
Appendix A Additional Tables
@@
Appendix B Game Library Design
@@
Table 6: Competitive (Zero-sum) Game - Base
@@
Table 7: Competitive (Zero-sum) Game - High-stake
@@
Table 8: Competitive (Zero-sum) Game - Low-stake
@@
Table 9: Cooperation Games (Stag Hunt) - Base
@@
Table 10: Cooperation Games (Stag Hunt) - High-payoff
@@
Table 11: Cooperation Games (Stag Hunt) - Asymmetric-payoff
@@
Table 12: Mixed-Motive Games (Prisoner’s Dilemma) - Base
@@
Table 13: Mixed-Motive Games (Prisoner’s Dilemma) - High-punishment for Defection
@@
Table 14: Mixed-Motive Games (Prisoner’s Dilemma) - Low-punishment for Defection
@@
    Competitive (Zero-Sum) Games: These games model adversarial interactions where one player’s gain is exactly balanced by the other player’s loss. The game consists of three variations:
    • Baseline: A standard zero-sum game where players choose strategies to maximize their own payoff while minimizing their opponent’s.
@@
Table 16: Bayesian Coordination Games - Type 1 Matrix
@@
Table 17: Bayesian Coordination Games - Type 2 Matrix
@@
Table 18: Signaling Games - Sender Matrix
@@
Table 19: Signaling Games - Receiver Matrix
@@
Appendix C Prompt Design
@@
The first type involves the vanilla version of the prompts, with specific variations for the Sequential Game, Bayesian Coordination Game, and Signaling Game, as detailed in Appendix B. The second type includes zero-shot CoT prompts, which add two additional sentences to the base prompts.
@@
C.1 Vanilla Experiment Prompt
The prompts used for the baseline evaluation of LLMs are listed in Table 21, 22, 23, and 24. These prompts are repeatedly sent using the LLMs’ API or cloud services’ API.
C.2 Demographic Feature Embedding
@@
C.3 Chain-of-Thought Embedding

### v3

@@ Appendix B Additional Results
Table 7: Strategic Reasoning Performance ($\tau$) Across Different Game Settings Include Distilled Models
@@
Table 8: Summary Statistics of Leading Models’ Completion Token Numbers for Baseline Games
Competitive games Cooperative games Mixed-strategy games
GPT-o1 GPT-o3-mini DeepSeek-R1 GPT-o1 GPT-o3-mini DeepSeek-R1 GPT-o1 GPT-o3-mini DeepSeek-R1
mean 7747.27 8938.33 10979.93 3197.80 5966.80 1764.67 3472.87 5944.07 2318.33
min 3994 6200 9050 2431 4786 1001 2585 4586 1599
max 14858 13223 13350 5690 7100 2931 5020 7659 3742
Table 9: Strategic Reasoning Performance ($\tau$) Across Different Game Settings in CoT
@@
Table 10: Context Controller ($\gamma$) Across Different Game Settings
Complete Information Game Incomplete Information Game
Competitive
Games
Cooperation
Games
Mixed-motive
Games
Seq.
Games
Bayesian
Games
Signaling
Games
BL HS LS BL HP AP BL H-Pun L-Pun p=0.5 p=0.9
GPT-4o 1.486 0.595 4.076 0.000 0.000 5.000 0.846 0.950 0.236 0.000 0.000 0.000 34.198
GPT-o1 0.064 0.029 1.231 5.000 0.047 5.000 5.001 5.276 0.395 10.152 1.000 0.000 0.676
GPT-o3-mini 1.482 0.595 1.316 5.000 0.041 5.000 5.001 5.276 0.120 0.089 5.000 0.071 35.734
Gemma-V2 22.898 0.592 3.010 0.000 0.000 1.045 0.031 0.317 0.129 0.000 0.000 0.000 52.362
Gemini-2.0 1.471 0.568 3.119 1.004 0.023 5.000 5.001 5.276 0.300 0.044 1.000 0.000 0.000
Claude-3-Opus 0.000 0.003 17.497 0.000 0.061 5.000 5.000 5.276 0.355 0.000 0.000 0.000 38.675
Granite-3.1-8B 0.125 5.000 0.144 0.000 0.000 5.000 0.000 0.000 0.000 0.000 5.000 0.000 0.000
Granite-3.1-MoE 2.383 0.931 3.874 5.000 0.040 5.000 5.000 1.010 2.860 0.065 5.000 0.054 0.000
LLaMA-3.1-405B 1.461 0.589 16.412 0.000 0.000 5.000 0.025 0.441 0.000 0.000 0.000 0.000 37.102
LLaMA-3.1-8B 0.000 0.000 0.000 1.009 0.000 0.000 0.337 5.861 0.081 0.000 0.000 0.000 48.384
InternLM-V2 16.153 0.000 0.000 0.000 0.000 5.000 0.000 0.593 0.061 0.000 0.000 0.000 42.306
QwQ-32B 0.045 0.921 0.143 5.000 0.032 0.056 6.609 0.842 0.882 0.109 5.000 0.331 48.972
GLM-4 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 38.796
DeepSeek-V2.5 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 41.697
DeepSeek-V3 0.706 0.569 21.732 5.000 0.041 5.000 5.001 5.276 0.120 0.054 5.000 0.078 52.656
DeepSeek-R1 13.719 0.583 19.847 0.000 0.000 1.045 0.000 0.014 0.156 0.000 0.000 0.000 40.615
Note: Abbreviations: HS = High Stake, LS = Low Stake, BL = Baseline, HP = High Payoff, AP = Asymmetric Payoff, H-Pun = High Punishment, L-Pun = Low Punishment.
@@
Table 13: Likelihood Across Different Game Settings
Complete Information Game Incomplete Information Game
Competitive
Games
Cooperation
Games
Mixed-motive
Games
Seq.
Games
Bayesian
Games
Signaling
Games
BL HS LS BL HP AP BL H-Pun L-Pun p=0.5 p=0.9
GPT-4o -1.498 -1.363 -1.932 -1.386 -1.386 -1.368 -0.635 -0.646 -1.129 -0.693 -1.386 -1.386 -1.045
GPT-o1 -1.849 -2.003 -1.020 -1.086 -1.304 -1.160 -0.023 -0.023 -0.801 -1.368 -1.384 -1.386 -0.742
GPT-o3-mini -1.427 -1.066 -1.169 -0.708 -1.253 -0.023 -0.023 -0.023 -1.336 -1.256 -0.708 -1.332 -0.780
Gemma-V2 -2.045 -1.484 -1.818 -1.386 -1.386 -1.386 -1.375 -1.379 -1.326 -1.386 -1.386 -1.386 -0.675
Gemini-2.0 -1.603 -1.663 -1.623 -1.384 -1.370 -0.650 -0.023 -0.023 -0.965 -1.365 -1.366 -1.386 -1.099
Claude-3-Opus -2.197 -2.197 -2.194 -1.386 -1.177 -0.785 -0.490 -0.023 -0.934 -1.386 -1.386 -1.386 -0.657
Granite-3.1-8B -2.164 -2.197 -2.135 -1.386 -1.386 -1.372 -1.386 -1.386 -1.386 -1.386 -1.370 -1.386 -2.197
Granite-3.1-MoE -1.945 -1.897 -1.901 -0.861 -1.280 -0.367 -0.454 -0.478 -0.303 -1.162 -0.752 -1.364 -2.197
LLaMA-3.1-405B -1.670 -1.819 -2.154 -1.386 -1.386 -1.386 -1.381 -1.373 -1.386 -1.386 -1.386 -1.386 -0.693
LLaMA-3.1-8B -2.197 -2.197 -2.197 -1.386 -1.386 -1.386 -1.367 -1.382 -1.373 -1.386 -1.386 -1.386 -0.996
InternLM-V2 -2.164 -2.197 -2.197 -1.386 -1.386 -1.368 -1.386 -1.339 -1.386 -1.386 -1.386 -1.386 -0.903
QwQ-32B -2.162 -1.884 -2.104 -1.194 -1.359 -0.841 -1.273 -0.987 -0.638 -1.281 -1.330 -1.216 -1.080
GLM-4 -2.197 -2.197 -2.197 -1.386 -1.386 -1.386 -1.386 -1.386 -1.386 -1.386 -1.386 -1.386 -0.871
DeepSeek-V2.5 -2.197 -2.197 -2.197 -1.386 -1.386 -1.386 -1.386 -1.386 -1.386 -1.386 -1.386 -1.386 -1.078
DeepSeek-V3 -2.195 -1.683 -2.104 -0.708 -1.253 -0.023 -0.023 -0.023 -1.336 -1.349 -1.086 -1.323 -0.876
DeepSeek-R1 -2.148 -2.111 -2.193 -1.386 -1.386 -1.386 -1.386 -1.385 -1.277 -1.386 -1.386 -1.386 -0.758
Note: Abbreviations: HS = High Stake, LS = Low Stake, BL = Baseline, HP = High Payoff, AP = Asymmetric Payoff, H-Pun = High Punishment, L-Pun = Low Punishment.
Appendix C Game Library Design
@@
Table 14: Competitive Games
(a) Base
@@
(b) High-stake
@@
(c) Low-stake
@@
Table 15: Cooperation Games (Stag Hunt)
(a) Base
@@
(b) High-payoff
@@
(c) Asymmetric-payoff
@@
Table 16: Mixed-Motive Games (Prisoner’s Dilemma)
(a) Base
@@
(b) High-punishment
@@
(c) Low-punishment
@@
    Competitive Games: These games model adversarial interactions where one player’s gain is exactly balanced by the other player’s loss. The game consists of three variations:
    • Baseline: A standard competitive game where players choose strategies to maximize their own payoff while minimizing their opponent’s.
@@
Table 18: Incomplete Information Games Payoff Matrices
(a) Bayesian Coordination Games
(b) Type 1
@@
(c) Type 2
@@
(d) Signaling Games
(e) Sender Payoffs
@@
(f) Receiver Payoffs
@@
Appendix D Prompt Design
@@
The first type involves the vanilla version of the prompts, with specific variations for the Sequential Game, Bayesian Coordination Game, and Signaling Game, as detailed in Appendix C. The second type includes zero-shot CoT prompts, which add two additional sentences to the base prompts.
@@
D.1 Vanilla Experiment Prompt
The prompts used for the baseline evaluation of LLMs are listed in Table 20, 21, 22, and 23. These prompts are repeatedly sent using the LLMs’ API or cloud services’ API.
D.2 Demographic Feature Embedding
@@
D.3 Chain-of-Thought Embedding